On January 2, 2019, a parent sells a building with original cost of $100,000 and accumulated depreciation of $25,000 to its wholly-owned subsidiary for $60,000. The estimated remaining life of the building is 5 years, and straight-line depreciation is appropriate. On the December 31, 2021, the subsidiary still owns the building.

The net effect of the working paper eliminations (I) for 2021 for this intercompany building sale:

A.

Increase accumulated depreciation by $34,000

B.

Increase investment in subsidiary by $6,000

C.

Add $15,000 to the original cost of the building

D.

Reduce 2021 depreciation expense by $3,000
Answer #1

Answer:- D. Reduce 2021 depreciation expense by $3,000

Explanation:-

Depreciation charged = (100,000 -25,000) / 5 = 15,000

Depreciation in 2021 = 60,000 / 5 = $12,000

So, the difference amount of $3,000 should be eliminated by Reducing the depreciation expense.

Please give a Thumbs up ?. Thanks!!